A regular hexagon of side 10 root 3 m is kept at horizontal surface . A particle is projected with velocity v m/s at an angle b from horizontal surface such that it will just touch the all four corners of regular hexagon.


Then answer the following


1. The maximum height from horizontal surface attained by projectile.


Ans. - 35m


2. The velocity of projectile at maximum height.


Ans.     5 root 3 m/s


 

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear anurag

let projectile is projected at a distance a from the corner of the base so its range must be

 R = a+10√3 + a

      = 2a +10√3

so 2a + 10√3 = V2 sin2b/2g 

 or  g/v2cos2b = sinb/(2a+10√3)  ....................1

select first two corner point of hexagon ,if the projectile pass through these point then by symmetry it will also pass through oteher two ppoints

 points are ( a-10√3cos60 , 10√3sin60)   and  (a,20√3sin60)

use the general equation

y =x tanb - gx2/2v2cos2b

put above these two point and put the value of g/v2cos2b = sinb/(2a+10√3)

then u will get 2 equation in a and tanb 

solve it  and then easly u can get desired result

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin